Two buses leave Boston at the same time traveling in opposite directions. One bus travels at 53 mph and the other at 64 mph. How soon will they be 936 miles apart?
Answer by fractalier(6550)

Since they cover a total of (53+64) = 117 miles per hour.
Now time = distance over rate or
t = d/r = 936/117 = 8 hours